OpenStudy (anonymous):

Please if anyone can help me with this, i;m trying to learn how to do this type of a problem, solve for x using logs 10^(x+8) = 6e^(3-x) please show me how to solve this

jimthompson5910 (jim_thompson5910):

take the natural log of both sides 10^(x+8) = 6e^(3-x) ln[ 10^(x+8) ] = ln[ 6e^(3-x) ] ln[ 10^(x+8) ] = ln(6) + ln( e^(3-x) ) (x+8)*ln(10) = ln(6) + (3-x)*ln( e ) (x+8)*ln(10) = ln(6) + (3-x)*1 (x+8)*ln(10) = ln(6) + 3-x do you see how to finish up?

OpenStudy (anonymous):

@jim_thompson5910, to be honest, not a clue how to finish it. keep coming up with the wrong answer

jimthompson5910 (jim_thompson5910):

(x+8)*ln(10) = ln(6) + 3-x x*ln(10)+8*ln(10) = ln(6) + 3-x x*ln(10)+8*ln(10)+x = ln(6) + 3 x*ln(10)+x = ln(6) + 3 - 8*ln(10) x(ln(10)+1) = ln(6) + 3 - 8*ln(10) x = ( ln(6) + 3 - 8*ln(10) )/( ln(10)+1 ) So the exact answer is \[\Large x = \frac{\ln(6) + 3 - 8\ln(10)}{\ln(10)+1}\]

OpenStudy (anonymous):

awesome my brother figured it out

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

thank you so much!